# On Nanocones as a Gravitational Analog System

### v1

@@ IV Universal Force Field in Molecular Mechanics
The complexity of multi-electron atoms makes it challenging to apply quantum mechanics directly to solve their problems. In such scenarios, we must resort to approximate methods alongside the many-electron wave function trial $\psi\left(\vec{r_{1}},\vec{r_{2}},\vec{r_{3}},....,\vec{r_{n}}\right)$ bretonnet2017basics ; viana2004quantitative to solve the appropriated Schrodinger equation, and therein obtain the behavior of these systems. Due the computation limitations, molecular mechanics, a classical approach, becomes indispensable for large atomic systems, aiding in understanding molecular properties and behavior in diverse settings. At the core of molecular mechanics lies the Universal Force Field (UFF), a framework crucial for simulating molecular interactions rappe1992uff ; jaillet2017uff .
The UFF is established empirically, parameterizing potential energy functions based on data mackerell2004empirical ; bradley1972ap . It simplifies by focusing solely on nuclear positions, neglecting electron movements, and computes an effective potential through the sum of atomic contributions potentials tajkhorshid2002control ; leach2001molecular , i.e., treats the atoms as particles. The force field encompasses various atomic interactions, including bond stretching, bending, rotations, cross-interactions, and non-bonded interactions. The bonded potential $V_{bonded}$ comprises terms for bonded pairs, valence angles, bending, consecutive stretching, consecutive bending, and cross-interactions. Conversely, $V_{non\,bonded}$ captures non-bonded interactions between atom pairs.
The UFF expression (42) combines these terms into a comprehensive potential energy function, while the torsional potential $V_{torsion}$ (43) handles torsion terms of the combined atoms. These formulations enable detailed simulations of molecular dynamics, essential for exploring the behavior of complex systems. The Universal Force Field is comprised of two general terms given by
$V_{\vec{r_{N}}}=V_{bonded}+V_{no\,bonded}$ (40)
for a system with N nuclei. The potential $V_{bonded}$ is given by
$V_{bonded}$ $=\sum_{d}V(d)+\sum_{\theta}V(\theta)+\sum_{\phi}V(\phi)+\sum_{\chi}V(\chi)$
$+\sum_{b,b^{{}^{\prime}}}V(b,b^{{}^{\prime}})+\sum_{\theta,\theta^{{}^{\prime}}}V(\theta,\theta^{{}^{\prime}})+\sum_{b,\theta^{{}^{\prime}}}V(b,\theta^{{}^{\prime}})$
$+\sum_{\theta,\theta^{{}^{\prime}}}V^{{}^{\prime}}(\theta,\theta^{{}^{\prime}})+\sum_{{\chi},\chi^{{}^{\prime}}}V(\chi,\chi^{{}^{\prime}})\,.$ (41)
The potential $V_{no\,bonded}$ is
$\sum_{i=1}^{N}\sum_{j=i+1}^{N}[V_{Coulomb}+V_{Lennard-Jones}]\,.$
Finally, the Universal Force Field is described by the potential
$V_{\vec{r_{N}}}$ $=\sum_{b}D_{b}\big{[}1-\exp^{-a(b-b_{0})}\big{]}^{2}+\sum_{\theta}H_{\theta}(\theta-\theta_{0})^{2}$
$+\sum_{\theta}H_{\theta}(\theta-\theta_{0})^{2}+\sum_{b,b^{{}^{\prime}}}F_{bb^{{}^{\prime}}}(b-b_{0})(b^{{}^{\prime}}-b_{0}^{{}^{\prime}})$
$+\sum_{\theta,\theta^{{}^{\prime}}}F_{\theta\theta^{{}^{\prime}}}(\theta-\theta_{0})(\theta^{{}^{\prime}}-\theta_{0}^{{}^{\prime}})+\sum_{b,\theta}F_{b\theta}(b-b_{0})(\theta-\theta_{0})$
$+\sum_{i=1}\sum_{j=i+1}\Bigg{(}4\epsilon_{ij}\Bigg{[}\Bigg{(}\frac{\sigma_{ij}}{r_{ij}}\Bigg{)}^{12}-\Bigg{(}\frac{\sigma_{ij}}{r_{ij}}\Bigg{)}^{6}\Bigg{]}+\frac{q_{i}q_{j}}{2\pi\epsilon_{0}r_{ij}}\Bigg{)}$
$+V_{torsion}\,.$ (42)
The first sum in the UFF potential function pertains to bonded atomic pairs, accounting for the interactions between atoms directly linked by chemical bonds. The second term encompasses all valence angles within the molecule, defined by the angle between three atoms A-B-C where A and C are bonded to B. This angle follows a harmonic potential model, reflecting the natural tendency of bonded atoms to maintain specific angles. Moving on, the third term addresses bending interactions, capturing the propensity of atoms to deviate from their ideal bond angles. The fourth term sums up consecutive stretching terms, representing the energy associated with stretching bonds between pairs of atoms. Similarly, the fifth term focuses on consecutive bending interactions, highlighting the energy changes when atoms deviate from their ideal bond angles successively. The third to sixth terms encompass cross-interaction phenomena, such as cross-terminus, out-of-plane tension, quadratic stretching, quadratic bond-angle interaction, quadratic angle-angle interaction, and out-of-plane tension interaction. These terms capture complex molecular behaviors arising from combinations of stretching, bending, and torsional motions, as well as interactions involving multiple bonds and angles. Finally, the last term accounts for non-bonded interactions, considering all pairs of atoms (i and j) in the molecule. This term encapsulates electrostatic interactions (Coulombic) and van der Waals forces (Lennard-Jones), among others, which contribute to the overall energy of the molecular system leach2001molecular .
The torsional potential, denoted as $V_{torsion}$, is expressed as
$V_{torsion}$ $=\sum_{\phi}H_{\phi}\big{[}1+s\,cos(n\phi)\big{]}$
$+\sum_{\chi}H_{\chi}\chi^{2}+\sum_{\theta,\theta^{{}^{\prime}}}F_{\theta\theta^{{}^{\prime}}\phi}(\theta-\theta_{0})(\theta^{{}^{\prime}}-\theta_{0}^{{}^{\prime}}cos(\phi))$
$+\sum_{\chi,\chi^{{}^{\prime}}}F_{\chi\chi^{{}^{\prime}}}\chi\chi^{{}^{\prime}}\,.$ (43)
In equation (43) first term represents the sum of all torsion terms. The second term represents the proper and improper torsion terms (symmetrical and anti-symmetrical). The third term represents 2 bending and a torsion, and the fourth term represents the sum of the terms that include 1 bending and a torsion.
The development of UFF was a significant leap forward in creating a versatile and comprehensive force field because it uses a combination of experimental data and high-level quantum mechanical calculations to parameterise the force field for different types of atoms and their interactions mayne2013rapid ; vanommeslaeghe2014molecular . One of UFF’s major features is the use of atomic hybridization states to refine the parameterization process, which allows the force field to adapt to different bonding environments within the same element.
@@
The versatility of UFF makes it suitable for a wide range of applications in molecular modeling. This has significant implications for fields such as materials science, biochemistry and pharmacology mackerell2004empirical . For example, in drug design, UFF can be used to model the interaction between a drug molecule and its target protein, providing information on the binding affinities and conformational changes involved tute1995drug . In materials science, UFF facilitates the study of molecular structures and properties of polymers, nanomaterials and complex alloys. By accurately modeling the forces within these materials, researchers can predict mechanical properties, stability and reactivity, guiding the design of new materials with customized properties.
The wide applicability and parameterization of the Universal Force Field for almost all elements are among its greatest strengths. This universality makes it a valuable tool for researchers working with diverse chemical systems without the need to switch between different force fields. However, this generality also introduces certain limitations. As the UFF is widely applicable, it may not have the precision of specialized force fields tailored to specific types of molecules or interactions.
In our molecular mechanics calculations for this study, we utilized the Universal Force Field (UFF) within the computational framework provided by the Forcite code rappe1992uff . This approach employs classical mechanics, adhering to Newtonian principles with an empirical potential function to describe molecular interactions. The Forcite code executes calculations by iteratively minimizing the potential energy, a process crucial for achieving a stable system. The iterative procedure aims to reach the global minimum of the potential energy surface, indicating a state where the forces on the system balance out to zero. This self-consistency cycle persists until the convergence criteria are met, typically after multiple iterations. In order to ensure accuracy, we utilized an integration grid to finely sample the potential energy landscape during geometric optimizations. Geometric optimization involves adjusting the atomic coordinates iteratively to locate the configuration that corresponds to the minimum energy state. Our convergence criteria included limiting the maximum atomic displacement to $5.0\times 10^{-3}$ angstroms, the maximum energy change to $1\times 10^{-4}$ kcal/mol, the maximum force to $1.0\times 10^{-3}$ Hartree per angstrom, and the maximum stress to 0.015 gigapascals. These stringent criteria were essential for achieving precise and reliable results in our calculations, ensuring that the molecular configurations obtained were energetically favorable and physically meaningful.
@@ V Results and analysis
The total energy of the structure has several contributions, such as those associated with the terms in equation (43). Part of this energy, i.e., the one associated with the torsion potential, has geometric origins and it is global, meaning it is linked to the overall shape of the structure. This energy is caused by the disclination defect and is not expected to be significantly affected by other atomic phenomena, such as the number of atoms in the structure. Given that the potential from the Universal Force Field (UFF) is empirical, we can regard its results as quasi-experimental. Our goal is to identify this “energy derived of experimental data” with the torsion energy described by TEGR.
In this section, we consider disclination defects in graphene sheets in subsection V.1 and in boron nitride sheets, graphene-like sheets with boron and nitrogen atoms alternating positions, in subsection V.2. Energy will be measured in electron volts (eV) and distances in angstroms ($\AA$ = $10^{-10}$ meters) unless otherwise stated.
@@ V.1 Graphene nanocones
The second step is to construct the Volterra process for the graphene sheet. We consider distinct sheets of various radii that are nearly circular. Given the hexagonal symmetry, it is impossible to achieve a perfectly circular sheet, so we anticipate another minor error due to the deviation from a perfect circle of radius $R$ in the integration contained in (39). To construct the nanocone, we cut a sheet, as demonstrated for Sheet 1 (with radius $2R_{1}=48.765\AA$), for an angle of $60^{\circ}$, as shown in Figure 1.
Figure 1: Graphiccal representation of the removal of a section $\Delta\phi=60^{\circ}$ from graphene Sheet 1.
@@
Figure 3: Energy associated with the torsion potential (in eV) for the nanocones as a function of the disclination angle deficit $\Delta\phi$. The dots indicate the data from the simulation, and the lines represent the linear fit.
@@
In order to strengthen our results, we considered three additional sheets, namely Sheet 2 (with radius $54.5602R_{2}=$ 54.560 Å), Sheet 3 (with radius $68.8862R_{3}=$ 68.886 Å), and Sheet 4 (with radius $98.3802R_{4}=$ 98.380 Å), and performed the same procedure, i.e., four nanocones for each sheet. Thus, a total of 16 simulations were conducted for graphene sheets. The results can also be seen in Figure 3. The linear fits are
@@
for Sheets 2, 3, and 4, respectively. The correlations for the linear fits are $0.996$, $0.998$, and $0.991$, respectively.
TEGR predicts an energy independent of the radius, characterized by a universal constant (specific to the material) multiplied by the disclination angle. This universal constant likely depends on an unknown coupling constant $k$ and the thickness of the sheet $L$. Since we do not know the exact value of constant $k$, we can only rely on qualitative corroboration for equation (39). Further discussion on the nature of this constant in graphene and BN will be presented in Section VI.
@@
V.2 Boron nitrite nanocones
@@
We generated boron nitride (BN) sheets, known for their structural and thermal properties similar to graphene aldalbahi2015variations . Boron nitride possesses insulating properties and exhibits a white coloration. Unlike graphene, boron nitride sheets maintain strength with increasing layers and boast a high Young’s modulus ($0.865TPa$) falin2017mechanical . These characteristics make boron nitride suitable for forming stable nanocones.
@@
$E_{6BN}$ $=9.226\,\Delta\phi+0.164\,,$ (54)
along with their respective correlations: $0.992$, $0.995$, $0.996$, $0.996$, $0.981$, and $0.950$ for Sheets 1 to 6, respectively. All correlations are close to unity.
In equations (49-54), we also note a small constant term, albeit with a significantly smaller magnitude. Furthermore, we observe a linear increase in the coefficient $\alpha$ with the radius of the structure but, like for graphene, not in a linear manner. Sheet BN 1 had a radius less than half of Sheet BN 6, yet this difference did not affect the ratio of the linear coefficients in the same magnitude.

### v2

@@ IV Universal Force Field in Molecular Mechanics
The complexity of multi-electron atoms makes it challenging to apply quantum mechanics directly to solve their problems. In such scenarios, we must resort to approximate methods alongside the many-electron wave function trial $\psi\left(\vec{r_{1}},\vec{r_{2}},\vec{r_{3}},...,\vec{r_{n}}\right)$ bretonnet2017basics ; viana2004quantitative to solve the appropriated Schrödinger equation, and therein obtain the behavior of these systems. Due to the computational limitations, molecular mechanics, a classical approach, becomes indispensable for large atomic systems, aiding in understanding molecular properties and behavior in diverse settings. At the core of molecular mechanics lies the UFF, a framework crucial for simulating molecular interactions rappe1992uff ; jaillet2017uff . The UFF is established empirically, parameterizing potential energy functions based on experimental data mackerell2004empirical ; bradley1972ap .
It makes things easier by only looking at the positions of the nuclei and ignoring the movements of the electrons. It then figures out the effective potential by adding up the potentials of all the atoms, which means it treats the atoms like particles. A schematic representation illustrating each type of interaction is shown in Figure 1
Figure 1: A geometrical representation of some interactions described by the UFF potential is presented, highlighting some different types of interactions, such as the stretching, bending and twisting of atomic bonds, and non-bonded interactions.
@@
The development of UFF uses a combination of experimental data and high-level quantum mechanical calculations to parameterize the force field for different types of atoms and their interactions mayne2013rapid ; vanommeslaeghe2014molecular . The primary objective of UFF is to achieve molecular equilibrium geometries and ascertain material properties at a reduced computational cost. The versatility of UFF makes it suitable for a wide range of applications in molecular modeling. This has significant implications for fields such as materials science, biochemistry, and pharmacology mackerell2004empirical . For example, in drug design, UFF can be used to model the interaction between a drug molecule and its target protein, providing information on the binding affinities and conformational changes involved tute1995drug . In materials science, UFF facilitates the investigation of molecular structures and properties of polymers, nanomaterials, and complex alloys. By simply modeling the forces between atoms through an analytical form of potentials with some parameters, researchers not only can predict mechanical properties, stability, and reactivity but could guide the design of new materials with customized properties. The greatest strength of UFF is the wide applicability and parametrization for almost all elements. This universality makes it a valuable tool for researchers working with diverse chemical systems without the need to switch between different force fields.
In this methodology, the total potential energy of the system is represented as the superposition of multiple individual potential interactions, the general form utilized in obtaining the findings of this paper is shown at (40). Where the individual contributions to $V(R)$ describes the following components types in order: bond stretching, represented by a Morse potential; bond angle deformations; torsional (dihedral) interactions; improper (out-of-plane bending) interactions; couplings between distinct chemical bonds; couplings between bond angles; couplings between bond angles and bond lengths; interactions between torsional angles and bond angles; couplings between improper angles; the van der Waals potential (Lennard-Jones terms); and the electrostatic interactions between atomic pairs. The final double summation in  (40) refers to non-bonded interactions.
$V(R)$ $=\sum_{b}D_{b}\left[1-e^{-a(b-b_{0})}\right]^{2}+\sum_{\theta}H_{\theta}(\theta-\theta_{0})^{2}+\sum_{\phi}H_{\phi}\left[1-s\cos(n\phi)\right]$ (40)
$\quad+\sum_{\chi}H_{\chi}\chi^{2}+\sum_{b}\sum_{b^{\prime}}F_{bb^{\prime}}(b-b_{0})(b^{\prime}-b_{0}^{\prime})+\sum_{\theta}\sum_{\theta^{\prime}}F_{\theta\theta^{\prime}}(\theta-\theta_{0})(\theta^{\prime}-\theta_{0}^{\prime})$
$\quad+\sum_{b}\sum_{\theta}F_{b\theta}(b-b_{0})(\theta-\theta_{0})+\sum_{\phi}F_{\phi\theta\theta^{\prime}}\cos\phi(\theta-\theta_{0})(\theta^{\prime}-\theta_{0}^{\prime})$
$\quad+\sum_{\chi}\sum_{\chi^{\prime}}F_{\chi\chi^{\prime}}\chi\chi^{\prime}+\sum_{i}\sum_{j>i}\left[\frac{A_{ij}}{r_{ij}^{12}}-\frac{B_{ij}}{r_{ij}^{6}}+\frac{q_{i}q_{j}}{r_{ij}\varepsilon}\right].$
In equation (40), $b$ represents the bond length between two atoms in a molecule, $\theta$ is the bond angle formed between three consecutive atoms, and $\phi$ is the torsional angle between four consecutively bonded atoms, defining the rotation about a bond. The term $\chi$ corresponds to the out-of-plane angle, which quantifies deviations from planarity in structures such as aromatic rings and carbonyl groups. The parameter $a$ controls the stiffness of the bond stretching energy in the Morse potential, while $\epsilon$ represents either the depth of the potential well in the Lennard-Jones interaction or the permittivity of the medium in the electrostatic term, depending on the context. The subscript 00 denotes equilibrium or reference values of structural parameters. Additionally, $b$ and $b^{\prime}$ refer to different bond lengths, and $V$ and $V^{\prime}$ indicate distinct potential energy surfaces.
The last term in equation (40) accounts for non-bonded interactions, considering all pairs of atoms (i and j) in the molecule and providing an energetic treatment of atomic interactions. In this term, $r_{ij}$ represents the distance between two interacting particles (nuclei), $q_{i}$ and $q_{j}$ denote the atomic charges, $\epsilon$ is the depth of the potential well (Lennard-Jones energy minimum), and $\sigma_{ij}$ is the characteristic distance at which the Lennard-Jones potential equals zero. This term encapsulates electrostatic interactions (Coulombic) and van der Waals forces (Lennard-Jones), among others, contributing to the overall energy of the molecular system leach2001molecular .
The parameters $D_{b}$, $H_{\theta}$, $H_{\phi}$, and $H_{\chi}$ determine the energetic contributions of different molecular deformations. The term $D_{b}$ represents the bond dissociation energy and appears in the Morse potential, describing the energy required to break a chemical bond. The constants $H_{\theta}$ and $H_{\phi}$ correspond to the force constants for bond angle bending and torsional rotation, respectively, governing the resistance of the molecule to angular deformations. The parameter $H_{\chi}$ is the out-of-plane bending force constant, which measures the resistance of planar structures, such as aromatic rings and carbonyl groups, to deviations from planarity. The terms $F_{bb^{\prime}}$, $F_{\theta\theta^{\prime}}$, $F_{b\theta}$, and $F_{\phi\theta\theta^{\prime}}$ represent cross-interaction parameters that account for couplings between different molecular distortions. Specifically, $F_{bb^{\prime}}$ and $F_{\theta\theta^{\prime}}$ describe interactions between pairs of bond lengths and bond angles, respectively, while $F_{b\theta}$ captures the coupling between bond stretching and angle bending. The term $F_{\phi\theta\theta^{\prime}}$ quantifies the influence of bond torsion on bond angle variations. Finally, the coefficients $A_{ij}$ and $B_{ij}$ appear in the Lennard-Jones potential, which describes van der Waals interactions between non-bonded atoms. The parameter $A_{ij}$ governs the strength of the repulsive interaction, typically associated with Pauli exclusion effects, while $B_{ij}$ determines the attractive interaction due to induced dipole-dipole forces. All these parameters are adjusted empirically or derived from quantum mechanical calculations to accurately reproduce the mechanical and structural properties of molecules.
Given the significance of energy torsion in this paper, we will subsequently emphasize just the terms that denote the torsional components of the potential energy. They are
$V_{torsion}$ $=\sum_{\phi}H_{\phi}\left[1-s\cos(n\phi)\right]$
$+\sum_{\phi}F_{\phi\theta\theta^{\prime}}\cos\phi(\theta-\theta_{0})(\theta^{\prime}-\theta_{0}^{\prime})$
$+\sum_{\chi}H_{\chi}\chi^{2}+\sum_{\chi}\sum_{\chi^{\prime}}F_{\chi\chi^{\prime}}\chi\chi^{\prime}\,.$ (41)
The rotation around a chemical bond; the coupling between dihedral angles and bond angles models how torsional motion interacts with bond bending; improper torsions (out-of-plane bending) account for deformations in planar systems such as aromatic rings and carbonyl groups; the coupling between different improper torsions represents interactions between multiple out-of-plane distortions. With that in mind, we conducted molecular mechanics calculations using the Forcite module of Materials Studiorappe1992uff . To achieve the equilibrium geometry, the program employs an iterative approach designed to attain the global minimum of the potential energy surface, signifying a state where the forces acting on the system are balanced to zero. This self-consistency cycle continues until the convergence conditions are satisfied, usually following several iterations. To assure precision, we employed an integration grid to sample the potential energy landscape during geometric optimizations. Geometric optimization entails the iterative modification of atomic coordinates to identify the configuration that represents the minimal energy state. Our convergence criteria encompassed restricting the maximum atomic displacement to $5.0\times 10^{-3}\AA$, the maximum energy variation to $4.34\times 10^{-7}eV$, the maximum force to $2.7\times 10^{-2}eV/\AA$ , and the maximum stress to $0.01\,GPa$. The tight criteria were crucial for obtaining accurate and dependable findings in our computations, guaranteeing that the resulting molecular configurations were energetically advantageous and physically significant. Subsequently, we calculated the aggregate of torsional energy components to evaluate the energy required for deforming the graphene surface on nanocones.
@@ V Results and analysis
The total energy of the structure has several contributions, including those related to (41). Part of this energy, i.e., the one associated with the torsion potential, has geometric origins and it is global, meaning it is linked to the overall shape of the structure. This energy is caused by the disclination defect and is not expected to be significantly affected by other atomic phenomena, such as the number of atoms in the structure. Given that the potential from the UFF is empirical, we can regard its results as quasi-experimental. Our goal is to identify this “energy derived from experimental data” with the torsion energy described by TEGR.
In this section, we consider disclination defects in graphene sheets in subsection V.1 and in boron nitride sheets, graphene-like sheets with boron and nitrogen atoms alternating positions, in subsection V.2. Energy will be measured in electron volts (eV) and distances in angstroms ($\AA$, $10^{-10}$ meters) unless otherwise stated.
@@ V.1 Graphene nanocones
The second step is to construct the Volterra process for the graphene sheet. We consider distinct sheets of various radii that are nearly circular. Given the hexagonal symmetry, it is impossible to achieve a perfectly circular sheet, so we anticipate another minor error due to the deviation from a perfect circle of radius $R$ in the integration contained in (39).
To determine the radius, we measure the distances between edge atoms connected by a straight line and select the largest distance as the structure’s diameter. The radius is then approximated as half of this diameter. This value is identified with the radius $R$ of the defect, as presented in equation (39) for the surface integral. To construct the nanocone, we cut a sheet, as demonstrated for Sheet 1 (with radius $2R_{1}=48.765\AA$), for an angle of $60^{\circ}$, as shown in Figure 2. The choice of this order of magnitude for the radius has no fundamental justification; it merely ensures a manageable size for constructing the structure and performing its computational simulation.
Figure 2: Graphical representation of the removal of a section $\Delta\phi=60^{\circ}$ from graphene Sheet 1.
@@
Figure 4: Energy associated with the torsion potential (in eV) for the graphene nanocones as a function of the disclination angle deficit $\Delta\phi$. The dots indicate the data from the simulation, and the lines represent the linear fit.
@@
To strengthen our results, we considered three additional sheets to mitigate potential coincidences that may arise from analyzing only a single system. In other words, we selected multiple distinct samples (sheets) to demonstrate the universality of the observed behavior. Another motivation for this approach is the absence of a perfect circle: smaller sheets deviate from the ideal circular shape assumed in (39), while larger ones tend to warp, thereby distorting the nanocones. By examining multiple sheets, we can observe the expected linear behavior both with and without significant deviations from a perfect circle and with or without pronounced nanocone warping. Consequently, we increased the radii as much as feasible while still allowing for the manual application of the Volterra process. The additional sheets are Sheet 2 (with radius $54.5602R_{2}=$ 54.560 Å), Sheet 3 (with radius $68.8862R_{3}=$ 68.886 Å), and Sheet 4 (with radius $98.3802R_{4}=$ 98.380 Å), and performed the same procedure, i.e., four nanocones for each sheet. Thus, a total of 16 simulations were conducted for graphene sheets. The results can also be seen in Figure 4. The linear fits are
@@
for Sheets 2, 3, and 4, respectively. The correlations for the linear fits are $0.996$, $0.998$, and $0.991$, respectively. The core result is not the value of the linear coefficients, but the strong linear correlation observed. All sheets exhibit correlations close to unity. It is expected that the linear coefficients will diverge, as the different sizes will influence the uncontrolled effects of circular deviation and warping of the structures.
TEGR predicts an energy independent of the radius, characterized by a universal constant (specific to the material) multiplied by the disclination angle. This universal constant likely depends on an unknown coupling constant $\kappa$ and the thickness of the sheet $L$. Since we do not know the exact value of constant $\kappa$, we can only rely on qualitative corroboration for equation (39). Further discussion on the nature of this constant in graphene and BN will be presented in Section VI.
@@
V.2 Boron nitride nanocones
@@
We generated boron nitride (BN) sheets, known for their structural and thermal properties similar to graphene aldalbahi2015variations . Boron nitride possesses insulating properties and exhibits a white coloration. Unlike graphene, boron nitride sheets maintain strength with increasing layers and boost a high Young’s modulus ($0.865TPa$) falin2017mechanical . These characteristics make boron nitride suitable for forming stable nanocones.
@@
$E_{6BN}$ $=9.640\,\Delta\phi+0.164\,,$ (52)
along with their respective correlations: $0.992$, $0.995$, $0.996$, $0.996$, $0.981$, and $0.993$ for Sheets 1 to 6, respectively. All correlations are close to unity.
In equations (47-52), we also note a small constant term, albeit with a significantly smaller magnitude. Furthermore, we observe a linear increase in the coefficient $\alpha$ with the radius of the structure but, like for graphene, not in a linear manner. Sheet BN 1 had a radius less than half of Sheet BN 6, yet this difference did not affect the ratio of the linear coefficients in the same magnitude.